A shop sells a single product for $1.50 per unit. The fixed monthly expenses are
$630 and the cost of each item is $0.60. How many items of the product must the
shop sell in this month to break even?
Thank you, Melissa.

Answer by venugopalramana(3286)

A shop sells a single product for $1.50 per unit. The fixed monthly expenses are
$630 and the cost of each item is $0.60. How many items of the product must the
shop sell in this month to break even?
LET THE NUMBER SOLD TO BREAK EVEN BE =N PER MONTH
THEIR COST OF PRODUCTION =FIXED COSTS +VARIABLE COSTS=630+0.6N
THEIR SALE VALUE =1.5N
AT BREAK EVEN.....S.V=C.P
1.5N=630+0.6N
1.5N-0.6N=0.9N=630
N=630/0.9=700